# Detecting Discrepancies Between AI-Generated and Natural Images Using Uncertainty

### v1

@@ 4.2 Results
Table 3: AI-generated image detection performance on GenImage. Except for WePe and RIGID, all methods require training on SD V1.4. The results of the baseline method are all from GenImage
Models
Methods Midjourney SD V1.5 ADM GLIDE Wukong VQDM BigGAN Average
ResNet-50 54.90 99.70 53.50 61.90 98.20 56.60 52.00 68.11
DeiT-S 55.60 99.80 49.80 58.10 98.90 56.90 53.50 67.51
Swin-T 62.10 99.80 49.80 67.60 99.10 62.30 57.60 71.19
CNNspot 52.80 95.90 50.10 39.80 78.60 53.40 46.80 58.63
Spec 52.00 99.20 49.70 49.80 94.80 55.60 49.80 64.41
F3Net 50.10 99.90 49.90 50.00 99.90 49.90 49.90 64.22
GramNet 54.20 99.10 50.30 54.60 98.90 50.80 51.70 65.66
DIRE 60.20 99.80 50.90 55.00 99.20 50.10 50.20 66.49
Ojha 73.20 84.00 55.20 76.90 75.60 56.90 80.30 71.73
LaRE 66.40 87.10 66.70 81.30 85.50 84.40 74.00 77.91
RIGID 81.54 68.72 72.35 84.15 68.57 78.98 93.02 78.19
WePe 79.17 75.57 76.07 79.20 79.00 90.60 89.27 81.27
Figure 4: Performance varies with variance
Figure 5: The effects of disturbed blocks.
Figure 6: Illustration of detection performance varying with perturbation intensity under different degradation mechanisms, (a): JPEG compression, (b): Gaussian blur, and (c): Gaussian noise.
4.3 Ablation study
In this section, we perform ablation experiments. Unless otherwise stated, experiments are conducted on ImageNet benchmark.
Robustness to Perturbations. Robustness to various perturbations is a critical metric for detecting generated images. In real-world scenarios, images frequently undergo perturbations that can impact detection performance. Following RIGID, we assess the robustness of detectors against three types of perturbations: JPEG compression (with quality parameter $q$)), Gaussian blur (with standard deviation $\sigma$), and Gaussian noise (with standard deviation $\sigma$). As illustrated in Figure 6, training-free methods generally exhibit superior robustness compared to training-based methods, with our method achieving the best overall performance.
The impact of the degree of perturbation. As shown in Figure 5, we explore the effect of the degree of perturbation on the performance of WePe. It can be seen that WePe is quite robust to the level of perturbation noise. It is only when the noise is very large or very small that it leads to a degradation in performance. When the noise level is small, the features obtained before and after the model perturbation are extremely similar, while when the noise level is very large, the features obtained before and after the model perturbation are extremely dissimilar, and these two cases will result in the inability to effectively differentiate between real and generated images.
Selecting which layers’ parameters to perturb? As shown in Figure 5, we explore the choice of which layers’ parameters to perturb would achieve good performance. The horizontal coordinates in the graph indicate that the first $k$ blocks are perturbed, not the $kth$ block. The experimental results exhibit that our method obtains good performance when the parameters of the first 9 to the first 20 blocks are chosen to be perturbed. This demonstrates the robustness of our method. In practice, we can select the layers to be perturbed by a small set of real and generated images. And when the generated images are not available, we can also use the probe to determine which layers are perturbed using only the real image. We describe our method in Appendix A.14.
The effect of models. In our experiments, we mainly used DINOv2 ViT-L/14 to extract features. We further explore the effect of using other models of DINOv2, including ViT-S/14, ViT-B/14, and ViT-g/14. In addition to this, we conduct experiments on the CLIP:ViT-L/14. As shown in Table 5, the performance on CLIP is not as good as on DINOv2. We hypothesize that the difference comes from the training approach of these models. CLIP learns features using image captions as supervision, which may make the features more focused on semantic information, whereas DINOv2 learns features only from images, which makes it more focused on the images themselves, and thus better able to capture subtle differences in real and generated images. When using DINOv2, our methods performs poorly when the capacity of the model is small. This may be due to the fact that larger models can better capture the differences between real and fake images. In our main experiments, to balance detection efficiency and detection performance, we use DINOv2 ViT-L/14.
The effect of perturbation type. In our experiments, we perturb the model parameters by adding Gaussian noise. We further explore other ways of perturbation, such as adding uniform noise or Laplace noise to the weight. In addition to this, we also explore the effect of MC Dropout, i.e., using dropout during inference. As shown in Table 5, all three weight perturbation methods achieve good performance, and outperform MC Dropout.
Table 4: The effect of models.
model      AUROC      AP DINOv2: ViT-S/14      72.83      71.63 DINOv2: ViT-B/14      81.82      80.64 DINOv2: ViT-L/14      87.99      85.04 DINOv2: ViT-g/14      84.92      81.83 CLIP: ViT-L/14      77.89      77.90
Table 5: The effect of type of perturbation.
5 Limitation
The proposed weight perturbation provides a simple and effective method for detecting generated images, yet we have not theoretically justified the widespread use of the method due to the inclusion of a variety of strong prior assumptions, such as the assumption about treating generated samples as OOD data. Therefore, our future work will focus on establishing the theoretical foundations of our method.
6 Conclusion
ETHIC STATEMENT
According to the Code of Ethics, our work does not raise any ethical concerns, as there are no human subjects, private datasets, harmful insights, or research integrity issues. Our work does not raise any ethical concerns. In particular, our work focuses on detecting AI-generated images to avoid the potential risk induced by the development of generative models.
REPRODUCIBILITY STATEMENT
To highlight the efforts that have been made to ensure reproducibility, we summarize materials facilitating reproducible results:
  • Theoretical results. Our work mainly focuses on the empirical investigation without explicit theatrical results.
  • Datasets. All involved datasets are publicly available. To facilitate reproducible results, we provide details in Sec 4 and Appendix A.15.
  • Open Source. We will release our code once the paper is accepted. If reviewers request, we will include an anonymized link to the code in our response.
Appendix A Appendix
A.1 Concerns about models being contaminated by generated images
With the proliferation of generated images, future large-scale models could become contaminated by such data, making it increasingly difficult to distinguish between natural and generated images. One potential solution is to employ machine unlearning (Yao et al., 2023) techniques to detect and address generated images. Machine unlearning focuses on removing the influence of specific data from pre-trained models, primarily due to privacy concerns. In this context, when generative images are incorporated into the training process of large-scale models, we can utilize machine unlearning techniques to eliminate the effects of these images on pre-trained models. This approach would help ensure that the features of natural and generated images remain distinct and separable.
A.2 Discussion on distribution discrepancy
In this paper, the core assumption we make is that there is data distribution discrepancy between natural and generated images. This assumption is valid for current generative models and has been confirmed by many works (Corvi et al., 2023; Tan et al., 2024; Ricker et al., 2024). This assumption is also the foundation of many generative image detection methods (we cannot distinguish between images that are indistinguishable).
Secondly, we observe that this discrepancy in data distribution can be captured by the representation space of a vision model pre-trained on a large number of natural images, i.e., there is feature distribution discrepancy between the generated and natural images, as shown in Figure 7. However, this remains an observation, and we have not found theoretical proof despite reviewing the literature. We only observe a similar phenomenon in UnivFD (Ojha et al., 2023), where the feature distribution discrepancy is observed in the representation space of CLIP:ViT-L/14.
That said, we can confirm the existence of feature distribution discrepancy of generated and natural images based on an important metric for evaluating generative models, the FID score. The FID score measures the feature distribution discrepancy between natural and generated images on the Inception network (Szegedy et al., 2015). When the FID score is 0, it indicates that the two distributions do not differ. However, even on these simple networks such as Inception v3, advanced generative models like ADM still achieve an FID score of 11.84, not to mention that on powerful models such as DINOv2, we observe significant feature distribution discrepancy.
A.3 Measuring feature distribution discrepancy with FID scores
We further use the ”FID” score to measure the difference in feature distribution between natural and generated images. To avoid the effects of categories, we compute the FID scores using the DINOv2 model on the LSUN-BEDROOM benchmark. For each category of images, we randomly select 5000 images for calculation. In addition to calculating the FID scores between natural images and generated images, we also calculate the FID scores between natural images and natural images. As shown in Table 6, the FID scores between natural images and generated images are significantly higher than the FID scores between natural images and natural images. Moreover, there is a clear positive correlation between the detection performance of WePe and the FID score. This result fully explains the existence of feature distribution discrepancy between natural and generated images on DINOv2, and demonstrates that WePe can effectively detect the feature distribution discrepancy.
FID score 1.09 18.25 52.31 59.94 80.44 116.06 130.00 132.75
AUROC 50.00 73.85 83.50 86.23 88.84 94.16 94.18 95.34
A.4 Comparison with Gradient Cuff
Gradient Cuff (Hu et al., 2024) focuses on detecting jailbreak attacks in Large Language Models (LLMs). It finds that the landscape of the refusal loss is more precipitous for malicious queries than for benign queries. And then it uses stochastic gradient estimation to estimate gradient and use the gradient norm as the decision score. Thus, we can also leverage this interesting work to identify the distribution discrepancy for detection. We use the similar way to estimate gradient as the decision score. As shown in Table 8, we surprisingly find that this method even surpasses WePe. This suggests that it is also possible to distinguish between natural and generated images by estimating the gradient.
A.5 WePe on large multi-modal models
In addition to CLIP, we further test the performance of WePe on BLIP (Li et al., 2022). As shown in Table 8, the performance of WePe is unsatisfactory on these multimodal models, which may be due to the fact that the image features of the multimodal models are more focused on semantic information, in line with our discussions.
A.6 Performance on adversarial examples
We further test WePe on adversarial examples. We simply add Gaussian noise (with different standard deviation $\sigma$) to the test samples to simulate the adversarial examples. We test three cases: adding noise on the natural image, adding noise on the generated image and adding noise on all images. As shown in Table 10, when noise is injected, the feature similarity between the clean model and the noisy model for the noisy image decreases, which leads to a change in the detection performance. To mitigate this effect, we can perform detection by perturbing the model multiple times and using the average similarity. The effect of noise is successfully mitigated by ensemble as shown in Table 10.
Table 7: Comparison with Gradient Cuff.
Method      AUROC      AP WePe      87.99      85.04 WePe + Gradient Cuff      89.36      90.62
Table 8: WePe on large multi-modal models.
Model      AUROC      AP DINOv2      87.99      85.04 CLIP      77.89      77.90 BLIP      68.25      64.68
Table 9: Performance on adversarial examples.
Case      AUROC      AP Clean images      87.99      85.04 Noisy natural images, $\sigma=0.05$      81.79      78.99 Noisy natural images, $\sigma=0.1$      76.16      73.16 Noisy natural images, $\sigma=0.15$      69.53      66.83 Noisy generated images, $\sigma=0.05$      88.03      85.82 Noisy generated images, $\sigma=0.1$      88.77      87.14 Noisy generated images, $\sigma=0.15$      90.25      88.99 Noisy natural and generated images, $\sigma=0.05$      84.94      81.98 Noisy natural and generated images, $\sigma=0.1$      81.46      78.25 Noisy natural and generated images, $\sigma=0.15$      78.19      75.32
Table 10: Mitigating the effects of image noise through ensemble.
Case      AUROC      AP Clean images      87.99      85.04 Noisy natural images, $\sigma=0.05,n=1$      81.79      78.99 Noisy natural images, $\sigma=0.05,n=5$      85.30      82.59 Noisy natural images, $\sigma=0.1,n=1$      76.16      73.16 Noisy natural images, $\sigma=0.1,n=5$      81.36      78.39 Noisy natural images, $\sigma=0.15,n=1$      69.53      66.83 Noisy natural images, $\sigma=0.15,n=5$      75.16      72.58
A.7 Concerns about hard samples
WePe relies on the model being pre-trained on a large dataset of natural images. Given the abundance of natural images, it is possible WePe may misclassify other natural images that are out-of-distribution, as AI-generated images, leading to false-negative errors. Thus, we follow previous work to leverage AUROC as the main metric the evaluate different methods. This is because AUROC is a pivotal metric for reflecting the false-positive and false-negative costs. The hard (natural) samples play a crucial role in detecting generated images. We will explore how to leverage hard samples to promote the detection performance in our future work.
A.8 Boosting performance on CLIP
As mentioned above, WePe does not perform satisfactorily on vision language models such as CLIP. Since CLIP needs to unite image features and text features, the image features need to be projected into the union space, which results in the projected image features being more focused on semantic information. For this reason, we can improve the performance of WePe on CLIP by using the features before projection for detection, as shown in Table 12.
Table 11: Boosting performance on CLIP.
Model      AUROC      AP DINOv2      87.99      85.04 CLIP      77.89      77.90 CLIP without projection matrix      84.82      84.20
Table 12: Comparison of detection times.
Method      Time (s) AEROBLADE      17.6 RIGID      3.7 WePe      4.5
A.9 Comparison of detection times.
Table 13: Evading detection of RIGID by adding noise to the generated image.
Model      AUROC      AP
RIGID      83.58      81.58
WePe      87.99      85.04
RIGID with noisy generated images      18.69      34.51
WePe with noisy generated images      88.77      87.14
Our method use a perturbed pre-trained model that is fixed during inferring all test samples. Thus, our method can be processed within two forward passes. This is equal to the cost of RIGID that requires two forward passes for clean and noisy images. However, RIGID can concatenate clean and noisy images in a mini batch and obtain detection results by with a single forward pass. AEROBLADE requires only one forward pass, but it needs to compute the reconstruction error of the image. This takes a long time to reconstruct at the pixel level. Besides, AEROBLADE needs to use a neural network to compute the LPIPS score, leading to computational complexity. As shown in Table 12, we compare the time required to detect 100 images under the same conditions. Since AEROBLADE needs to calculate the image reconstruction error, it has the lowest detection efficiency. RIGID can obtain detection results in a single forward pass by concatenating clean and noisy images, whereas WePe requires two forward passes, which results in WePe’s detection efficiency being inferior to RIGID’s. However, WePe can be parallelized across two devices to obtain the detection results in a single forward pass.
A.10 Comparison with RIGID
The main differences between WePe and RIGID are as follows:
  • The approach proposed by RIGID stems only from the phenomenon it observed: namely, that natural and generated images show different sensitivities to noise in the representation space of DINOv2. Instead, WePe explicitly proposes that there is distribution discrepancy between natural and generated images, and utilizes the difference in uncertainty to expose distribution discrepancy between natural and generated images.
  • RIGID utilizes the difference in sensitivity to noise between natural image and generated image for detection. Although the generated image is more sensitive, it is easy to think of a way to avoid detection, i.e., the user adds noise to the generated image and then submits it for detection. This approach easily removes the sensitivity of the generated image to noise. As shown in Table 13, after adding noise to a generated image, RIGID determines the noisy generated image as natural images. This is a fatal flaw of RIGID. However, WePe is not affected by this. WePe exposes the difference in distribution between the test image and the natural image through weight perturbation. Adding noise to the generated image further increases this difference, leading to even better detection performance.
A.11 Note on the unbiased assumption
In section 3.3, we make an assumption that the expected extracted feature by noised models is unbiased for that by the original model. We think this assumption is reasonable. Thanks to the over-parameterization of modern neural networks and advanced optimization algorithms (e.g., AdamW), it is a well-established fact that trained neural networks are usually smooth in the parameter space and show robustness to small weight perturbation (Novak et al., 2018). And this robustness is used in many applications, such as quantization (Gholami et al., 2022) and pruning (Liu et al., 2019). And in Figure 2, we also clearly show this robustness: when adding tiny noise to the model weights, the features remain almost unchanged.
A.12 SOFTWARE AND HARDWARE
We use python 3.8.16 and Pytorch 1.12.1, and seveal NVIDIA GeForce RTX-3090 GPU and NVIDIA GeForce RTX-4090 GPU.
@@ A.12 SOFTWARE AND HARDWARE
A.13 WePe with multiple perturbation
In our experiments, taking into account the detection efficiency, we perturb the model only once, and then calculate the feature similarity of the test samples on the clean and perturbed models. We further experiment with multiple perturbations and use the mean of the feature similarity of the test samples on the clean model and all the perturbed models as the criterion for determining whether or not the image is generated by the generative models. As shown in Figure 8,
A.14 Using natural images only to select which layers to perturb
In our experiments, we use a small set of natural images and generated images to pick the parameters that need to be perturbed. When all the generated images are not available, we can also use only the real images to select the layers that need to be perturbed. Specifically, we first perturb each block alone and calculate the similarity of the features on the model of the natural image before and after the perturbation, as shown in Table 14. We then sort the similarity and select the blocks with the highest similarity for perturbation. As shown in Table 15, selecting the parameters to be perturbed in this way also achieves good performance and has strong robustness.
@@ A.14 Using natural images only to select which layers to perturb
A.15 Details of Datasets
IMAGENET. The real images and generated images can be obtained at https://github.com/layer6ai-labs/dgm-eval. The images are provided by (Stein et al., 2023). The generative model includes: ADM, ADMG, BigGAN, DiT-XL-2, GigaGAN, LDM, StyleGAN-XL, RQ-Transformer and Mask-GIT. The resolution of real images and generated images are $256\times 256$. We crop the image randomly to $224\times 224$ resolution.
LSUN-BEDROOM. The real images and generated images can be obtained at https://github.com/layer6ai-labs/dgm-eval. The images are provided by (Stein et al., 2023). The generative model includes: ADM, DDPM, iDDPM, StyleGAN, Diffusion-Projected GAN, Projected GAN and Unleashing Transformers. The resolution of real images and generated images are $256\times 256$. We crop the image randomly to $224\times 224$ resolution.
GenImagThe real images and generated images can be obtained at https://github.com/GenImage-Dataset/GenImage. The images are provided by (Zhu et al., 2023). The generative model includes: Midjourney, SD V1.4, SD V1.5, ADM, GLIDE, Wukong, VQDM and BigGAN. The real images come from ImageNet, and different images have different resolutions. Following (Stein et al., 2023), we resize the image to $256\times 256$ resolution and adjust its format to keep the same with the generated images, then we randomly crop it to $224\times 224$ resolution to extract features. For the generated images, we report our processing in detail as follows:
  • Midjourney. The resolution of images generated by Midjourney is $1024\times 1024$, and we randomly crop them to $224\times 224$ resolution.
  • SD V1.4. The resolution of images generated by SD V1.4 is $512\times 512$, and we randomly crop them to $224\times 224$ resolution.
  • SD V1.5. The resolution of images generated by SD V1.5 is $512\times 512$, and we randomly crop them to $224\times 224$ resolution.
  • ADM. The resolution of images generated by SD V1.5 is $256\times 256$, and we randomly crop them to $224\times 224$ resolution.
  • GLIDE. The resolution of images generated by SD V1.5 is $256\times 256$, and we randomly crop them to $224\times 224$ resolution.
  • Wukong. The resolution of images generated by SD V1.5 is $512\times 512$, and we randomly crop them to $224\times 224$ resolution.
  • VQDM. The resolution of images generated by SD V1.5 is $256\times 256$, and we randomly crop them to $224\times 224$ resolution.
  • BigGAN. The resolution of images generated by SD V1.5 is $128\times 128$, and we fill them with zero pixels to $224\times 224$ resolution.

### v2

@@ Appendix H WePe on large multi-modal models
Table 11: AI-generated image detection performance (ACC, %) on DRCT-2M.
Method SD Variants Turbo Variants LCM Variants ControlNet Variants DR Variants Avg.
LDM SDv1.4 SDv1.5 SDv2 SDXL
SDXL-
Refiner
SD-
Turbo
SDXL-
Turbo
LCM-
SDv1.5
LCM-
SDXL
SDv1-
Ctrl
SDv2-
Ctrl
SDXL-
Ctrl
SDv1-
DR
SDv2-
DR
SDXL-
DR
CNNSpot 99.87 99.91 99.90 97.63 66.25 86.55 86.15 72.42 98.26 61.72 97.96 85.89 82.94 60.93 51.41 50.28 81.12
F3Net 99.85 99.78 99.79 88.60 55.85 87.37 63.29 63.66 97.39 54.98 97.98 72.39 81.99 65.42 50.39 50.27 71.13
CLIP/RN50 99.00 99.99 99.96 94.61 62.08 91.43 84.40 64.40 98.97 57.43 99.74 80.69 82.03 65.83 50.67 50.47 80.05
GramNet 99.40 99.01 98.84 95.30 62.63 80.68 71.19 69.32 93.05 57.02 89.97 75.55 82.68 51.23 50.01 50.08 76.62
De-fake 92.10 95.53 99.51 89.65 64.02 69.24 92.00 93.93 99.13 70.89 58.98 62.34 66.66 50.12 50.16 50.00 75.52
Conv-B 99.97 100.0 99.97 95.84 64.44 82.00 60.75 99.27 99.27 62.33 99.80 83.40 73.28 61.65 51.79 50.41 79.11
Ojha 98.30 96.22 96.33 93.83 91.01 93.91 86.38 85.92 90.44 89.99 90.41 81.06 89.06 51.96 51.03 50.46 83.46
DIRE 54.62 75.89 76.04 99.87 59.90 93.08 97.55 87.29 72.53 67.85 99.69 64.40 64.40 49.96 52.48 49.92 72.55
DRCT 94.45 94.35 94.24 95.05 96.41 95.38 94.81 94.48 91.66 95.54 93.86 93.50 93.54 84.34 83.20 67.61 91.35
FatFormer 96.52 95.31 93.27 91.99 92.87 91.78 88.15 87.48 92.82 91.76 90.28 86.99 88.19 65.92 60.15 55.13 85.53
WePe 92.38 67.18 65.88 74.05 75.62 72.23 66.82 62.46 66.88 77.25 75.41 74.92 80.34 63.98 59.65 59.68 70.92
WePe 97.06 96.03 94.76 96.45 96.59 97.81 93.54 92.66 96.29 94.43 96.69 96.17 95.72 75.99 73.32 69.78 91.45
Appendix N SOFTWARE AND HARDWARE
We use python 3.8.16 and Pytorch 1.12.1, and several NVIDIA GeForce RTX-3090 GPU and NVIDIA GeForce RTX-4090 GPU.
@@
Appendix O Using natural images only to select which layers to perturb
In our experiments, we use a small set of natural images and generated images to pick the parameters that need to be perturbed. When all the generated images are not available, we can also use only the natural images to select the layers that need to be perturbed. Specifically, we first perturb each block alone and calculate the similarity of the features on the model of the natural image before and after the perturbation, as shown in Table 12. We then sort the similarity and select the blocks with the highest similarity for perturbation. As shown in Table 13, selecting the parameters to be perturbed in this way also achieves good performance and has strong robustness.
@@
Appendix P Details of Datasets
IMAGENET. The natural images and generated images can be obtained at https://github.com/layer6ai-labs/dgm-eval. The images are provided by (Stein et al.,, 2023). The generative model includes: ADM, ADMG, BigGAN, DiT-XL-2, GigaGAN, LDM, StyleGAN-XL, RQ-Transformer and Mask-GIT.
LSUN-BEDROOM. The natural images and generated images can be obtained at https://github.com/layer6ai-labs/dgm-eval. The images are provided by (Stein et al.,, 2023). The generative model includes: ADM, DDPM, iDDPM, StyleGAN, Diffusion-Projected GAN, Projected GAN and Unleashing Transformers.
GenImage. The natural images and generated images can be obtained at https://github.com/GenImage-Dataset/GenImage. The images are provided by (Zhu et al.,, 2023). The generative model includes: Midjourney, SD V1.4, SD V1.5, ADM, GLIDE, Wukong, VQDM and BigGAN. The natural images come from ImageNet, and different images have different resolutions.
DRCT-2M. The natural images of DRCT-2M come from CoCo and can be obtained from https://cocodataset.org/#download. AI-generated images of DRCT-2M can be obtained from https://modelscope.cn/datasets/BokingChen/DRCT-2M/files, which are provided by (Chen et al.,, 2024). The generative model includes LDM, SDv1.4, SDv1.5, SDv2, SDXL, SDXL-Refiner, SD-Turbo, SDXL-Turbo, LCM-SDv1.5, LCM-SDXL, SDv1-Ctrl, SDv2-Ctrl, SDXL-Ctrl, SDv1-DR, SDv2-DR, SDXL-DR.
Appendix Q Implementation details
To balance detection performance and efficiency, we use DINOv2 ViT-L/14. We report the average results under five different random seeds and report the variance in Figure 10. In our experiments we find that perturbing the high layers may lead to a large corruption in the features of the natural images, resulting in sub-optimal results. Therefore, We do not perturb the high-level parameters. In DINOv2 ViT-L/14, the model has 24 transformer blocks, and we only perturb the parameters of the first 19 blocks with Gaussian perturbations of zero mean. The variance of Gaussian noise is proportional to the mean value of the parameters in each block, with the ratio set to 0.1. Considering the computational cost, we perturb the model only $1$ time, i.e., $n=1$. Multiple perturbations can further improve the performance as shown in Table 10. For WePe, we leverage LoRa (Hu et al.,, 2022) for parameter-effcient fine-tuning. The Lora layers are applied on the q_proj and v_proj layers of DINOv2. lora_rlora\_r and $lora\_\alpha$ are set to 8. And the model is optimized using the AdamW optimizer with a learning rate of $1\times 10^{-5}$, $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a weight decay of $0.01$. Following CNNspot (Wang et al.,, 2020), data augmentation techniques including JPEG compression and Gaussian blur are employed to enhance robustness. For the IMAGENNET and LSUN-BEDROOM benchmarks, the ProGAN dataset serves as the training set. For the GenImage benchmark, SDv1.4 dataset is used as training set. For the DRCT-2M benchmark, SDv2 dataset is used as training set. When testing, to ensure objectivity in calculating classification accuracy and mitigate biases arising from manually selected thresholds, following (Ojha et al.,, 2023), we automatically determine the optimal threshold by identifying the score that maximizes the separation between natural and generated images, based on their computed classification scores.